OpenStudy (anonymous):

I'm going to try to explain my question the best I can. Okay, I had to write an equation of the line perpendicular to the given line and containing the given point. I then had to write the answer in standard form. Here is the question: y=-5x+1; (10,0) I know the reciprocal is 1/5. So then I used point-slope formula to find an equation of the line. y-0=1/5(x-10) then I got y-0=1/5x-2 I then multiplied everything by 5 and got 5y-0=x-10. I then answered it X+5y=-10, but the book says X-5y=10. So my question is, when do you know to make it an addition/sub problem?

ganeshie8 (ganeshie8):

Your work is correct, till :- 5y-0=x-10

ganeshie8 (ganeshie8):

standard form is :- Ax + By = C

ganeshie8 (ganeshie8):

and usually, we want the x term to have positive coefficient

ganeshie8 (ganeshie8):

5y-0=x-10 so here, i wud simply subtract 5y both sides : 5y-0=x-10 -5y -5y 0 = x-5y - 10

ganeshie8 (ganeshie8):

next i wud add 10 both sides : 0 = x-5y - 10 +10 +10 10 = x - 5y its same as, x - 5y = 10

ganeshie8 (ganeshie8):

im not sure if that answers ur q :o

OpenStudy (anonymous):

OH okay, I get it now. I see what I did wrong. I need to pay more attention to the signs, thanks! I wanted to subtract 0 from -10, but that should of been adding zero to 10...

ganeshie8 (ganeshie8):

looks u nailed it :D

OpenStudy (anonymous):

^_^